OpenStudy (anonymous):

Simplifying radical expressions please help! Square root 20c^2 d^2

OpenStudy (anonymous):

\[\sqrt{20c ^{2} d ^{2}}\]

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

\[\sqrt{c^2}=c\]

OpenStudy (anonymous):

\[\sqrt{d^2}=d\]

OpenStudy (anonymous):

\[\sqrt{20c ^{6} d ^{5}}\]

OpenStudy (anonymous):

i put it in wrong

OpenStudy (anonymous):

\[\sqrt{20}=\sqrt{4\times 5}=\sqrt{4}\sqrt{5}=2\sqrt{5}\]

OpenStudy (anonymous):

@satellite73 will you look at my correction on the equation